write the slope intercept form of an equation for the line that passes through (-4,-3) and is paraell to thr graph of y = -3/7x-2
Answer by nerdybill(7384)

.
For a line to be parallel to:
y = -3/7x-2
BOTH lines MUST have the same slope
.
Notice:
y = -3/7x-2
is already in the "slope-intercept" form of:
y = mx + b
where
m is slope
b is y-intercept
.
So, we can see that our new line MUST have a slope of -3/7 AND crosses the point at (-4,-3) -- plug the two pieces of information into the "point-slope" form:
y - y1 = m(x-x1)
y - (-3) = (-3/7)(x-(-4))
y + 3 = (-3/7)(x+4)
multiplying both sides by 7:
7y + 21 = -3(x+4)
7y + 21 = -3x - 12
7y = -3x - 33
y = (-3/7)x - 33/7 (this is what they're looking for)